ABC Corp shares trade at $60 and you short 500 shares. The initial margin is 50% and your broker has a 35% maintenance margin.

(A) If the stock price rises to $65, what is your percentage margin?

(B) How far can the stock price rise before a margin call?
Answer #1

Value of 500 shares = 500*60 = 30000
Loan amount = 30000*50% = 15,000

A) Percentage Margin = (65*500 - 15000)/65*500 = 53.85%

b) Price to get margin call
35% = (500*P-15000)/(500P)
175P = 500P - 15000
P = 15000/(500-175) = 46.15

Price rise before margin call = 46.15

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well